H2O -> H2 + O2




A.How many moles of O2 are produced when 1.26 moles if H2O is reacted?




B. How many moles of H2O are needed to produce 55.7 moles of H2




C. If enough H2O is reacted to produce 3.40 moles of H2 then how many moles of O2 must have been made?

1
Expert's answer
2022-05-12T11:36:06-0400

2H2O -> 2H2 + O2

According to the equation, n (H2O) = n (H2) = 2 x n (O2)


A. n (H2O) = 2 x n (O2), n (O2) = 1/2 x (H2O) = 1/2 x 1.26 = 0.63 mol


B. n (H2O) = n (H2) = 55.7 moles


C. n (H2) = 2 x n (O2), n (O2) = 1/2 x (H2) = 1/2 x 3.4 = 1.7 moles